Q 11) A thread was tilted 5° from vertical and let go. The ball’s movement time from the maximum

deflection point is?

1) 0.1s

2) 0.5s

3) 1s

4) 2s


1
Expert's answer
2021-05-12T14:59:10-0400

"T=2\pi\sqrt{l/g}"


If "l=1\ m" then "T=2\cdot3.14\cdot\sqrt{1/9.8}=2\ (s)"


The time of movement of the ball from the maximum deflection point to the equilibrium position is


"2/4=0.5\ (s)"



2) "0.5\ (s)" . Answer